### v1

# COMET: Co-Optimization of a CNN Model using Efficient-Hardware OBC Techniques

Abstract
Convolutional Neural Networks (CNNs) are highly effective for computer vision and pattern recognition tasks; however, their computational intensity and reliance on hardware such as FPGAs pose challenges for deployment on low-power edge devices. In this work, we present COMET, a framework of CNN designs that employ efficient hardware offset-binary coding (OBC) techniques to enable co-optimization of performance and resource utilization. The approach formulates CNN inference with OBC representations of inputs (Scheme A) and weights (Scheme B) separately, enabling exploitation of bit-width asymmetry. The shift–accumulate operation is modified by incorporating the offset term with the pre-scaled bias. Leveraging inherent symmetries in Schemes A and B, we introduce four novel look-up table (LUT) techniques—parallel, shared, split, and hybrid—and analyze them to identify the most efficient options. Building on this foundation, we develop an OBC-based general matrix multiplication core using the im2col transformation, enabling efficient acceleration of a fixed-point modified LeNet-5 model. FPGA evaluations demonstrate that the proposed co-optimization approach significantly reduces resource utilization compared to state-of-the-art LeNet-5 based CNN designs, with minimal impact on accuracy.
Index Terms:
Convolutional Neural Networks (CNN), Field-Programmable Gate Array (FPGA), General Matrix-Multiply (GEMM), Look-up Table (LUT), Offset-Binary Coding (OBC).
I Introduction
Convolutional neural networks (CNNs) have become a cornerstone of modern deep learning, propelled by the rapid progress in artificial intelligence. Their ability to learn features directly from raw data in an end-to-end manner has enabled them to significantly surpass traditional handcrafted methods in key computer vision and pattern recognition tasks, including image classification, object detection, and semantic segmentation [1]. Owing to these advantages, CNNs have found widespread use in diverse domains such as industrial defect inspection, medical imaging, and others [2]. However, despite their success, conventional CNN inference relies on multiplication-intensive operations that are power-hungry and chip-area demanding, posing challenges for edge platforms. To address these challenges, CNNs have been increasingly implemented on FPGAs [3, 4, 5], which are particularly well-suited for CNN acceleration due to their inherent parallelism, reconfigurability, and efficient utilization of on-chip memory and DSP slices [6].
To optimize resource utilization on FPGA, many CNN accelerators adopt the im2col transformation [7] to restructure 2D convolutions into tile-based general matrix multiplication (GEMM) workloads, enabling the reuse of a shared multiply–accumulate (MAC) array across layers [8]. While this strategy improves computational throughput, it exacerbates the demand for DSP multipliers and increases pressure on on-chip memory bandwidth. To mitigate these constraints, various multiplication-reduction techniques have been explored, including Winograd minimal filtering [9], sparse convolution [10], and depthwise separable convolutions [11]. Although these methods effectively reduce the number of operations, they remain tied to DSP-centric architectures—limitations that are particularly pronounced in edge scenarios where both DSP availability and dynamic power budgets are highly constrained.
For efficient CNN implementation on resource-constrained hardware, asymmetric quantization is particularly useful when weight bit-width can be reduced while keeping input bitwidth fixed, albeit at the expense of accuracy [12]. To effectively evaluate such methods, it is essential to select representative models that balance computational complexity and accuracy. Lightweight architectures are particularly well-suited for exploring low-area, low-power design strategies, especially where minimizing energy consumption and silicon footprint is critical. While several works  [13, 14, 4] have targeted complex models such as MobileNetV2 and SqueezeNet—leveraging depthwise separable convolutions and compact network designs to optimize throughput—these approaches often depend on MAC operations implemented through DSP blocks. However, in severely resource-constrained environments, they can become inefficient due to their high computational and energy cost.
LeNet-5 model [15, 16] is often selected as a representative small-scale CNN because of its: (1) simple architecture and low parameter count, which make it suitable for resource-constrained hardware [17]; (2) well-established MNIST performance, providing a reproducible benchmarking baseline [18]; and (3) compatibility with operator types and problem scales typically targeted by GEMM-based frameworks [19]. Unlike the aforementioned models, it avoids architectural complexity, enabling focused studies of arithmetic-level optimizations and low-power, low-area trade-offs. Following this rationale, the present work adopts the classic LeNet-5 CNN as its evaluation benchmark. The model is compact enough to fit entirely on mid-range FPGAs, yet still encompasses representative convolutional, pooling, and fully connected layers. Recent studies [20, 21] consistently demonstrate that accuracy variations introduced by architectural modifications are largely independent of training factors. For instance, replacing the original tanh activations with ReLU leads to better performance in practice [20], while substituting large fully connected layers with global average pooling (GAP) offers a more efficient alternative [21].
Distributed arithmetic (DA) is a multiplier-less approach ideal for low-power, area-constrained hardware [22, 23]. It replaces MAC operations in inner-product computation (IPC) with precomputed look-up tables (LUTs) and a shift-accumulate (SA) unit. This involves representing one vector in two’s complement (TC) [24] or offset-binary coding (OBC) [25], and storing binary or offset binary combinations of the other vector in LUTs. While TC causes exponential LUT growth, OBC reduces this by half through symmetry. Recent works have optimized DA-based LSTMs using hardware-based LUTs [26, 27, 28, 29]. For instance, [26] proposed a parallel OBC-based LUT with pipelining, which reduced critical path delay (CPD) but increased hardware complexity. While [28] improved this with a more efficient approach based on TC. Furthermore, serial LUT designs using high-radix OBC and TC were proposed in [27, 29], incurring additional clock cycle latency. The serial-LUT suffers from limited reusability and significant delay for large dimensions, making pipelining essential but adding latency. To the best of our knowledge, only a few DA-specific implementations targeting either 1-point convolution or CNN acceleration (on ASICs) exist [30, 31, 32]; for example, [30] minimized the LUT size via input pairing, while [31] reduced memory usage through in-memory optimization.
Figure 1: Illustration of the im2col transformation of a 2D convolution into a GEMM operation, assuming no padding in the input feature map and stride=1.
Although various algorithmic optimizations have helped reduce MAC operation counts in CNNs, they still fundamentally rely on DSP-based multipliers, which remain a critical bottleneck for low-power edge deployments. Even advanced dataflow strategies [11, 3] that aim to maximize data reuse fail to eliminate this reliance on MAC units, thereby limiting the efficiency and scalability of accelerators targeting stringent energy and area constraints. This persistent challenge motivates the design of a fully DSP-free CNN architecture based on LeNet-5 model. In this work, we first adopt a streaming (im2col + tiled) strategy to partition the original feature map into multiple small matrix blocks and perform GEMM on each block, thereby avoiding large-scale IPCs in a single pass. To eliminate large pre-computed LUTs, we propose dynamic parallel LUT structure and its variant that enable flexible sharing and reuse via runtime computation of the partial sums. A CNN architecture relying on the OBC scheme is developed and integrated into a general-purpose GEMM core. Using im2col, standard CNN kernels are mapped and executed as OBC-accelerated GEMM operations. We propose COMET (Co-Optimization of a CNN Model using Efficient Hardware OBC Techniques), a framework for efficient CNN acceleration that leverages hardware-optimized OBC techniques to co-optimize performance and resource utilization. The main contributions of this work are as follows:
  • A compact im2col-based formulation of 2D CNN convolution layer using OBC representations of inputs (Scheme A) and weights (Scheme B) enables systematic analysis for hardware optimization.
  • As part of the COMET framework, four novel LUT architectures are introduced using efficient hardware OBC techniques. The shift–accumulate unit is enhanced to merge the pre-scaled bias and offset terms, enabling co-optimization of resource and energy efficiency across multiple clock frequencies.
  • LeNet-5 CNN kernels are mapped onto an OBC-GEMM framework with unified accumulation using Scheme A and Scheme B, and impact of quantization on the accuracy is also analyzed.
  • The design space of the proposed architectures is explored, implemented, and evaluated on an FPGA platform, providing a comprehensive analysis of design trade-offs such as resource utilization and energy efficiency.
The remainder of this paper is organized as follows. Section II introduces the OBC formulation of a standard CNN. Section III presents the proposed LUT architectures and analyzes their co-optimization trade-offs. Section IV discusses the design considerations and evaluation of the proposed OBC-GEMM module. Section V provides a comparative analysis of the proposed design against state-of-the-art CNN accelerators. Finally, Section VI concludes the paper.
Figure 2: OBC-based IPC for $K=4$ with (a) traditional LUT and SA-unit [25], (b) hardware LUT (w/o pipelining) and CSFA-based SA unit [26].
II CNN Formulation with OBC
As shown in Fig. 1, a standard CNN is a multi-layer pipeline in which each layer transforms input feature maps of size $H\times W$ into higher-level representations using convolutional kernels of size $K\times L$ and biases. The layer has $C$ input channels and $N$ output channels. Each output channel is obtained by convolving all $C$ input channels with their corresponding kernels via a sliding-window operation (A), summing across channels, and repeating this process for all $N$ outputs. For clarity, we consider the case of batch size $1$, although the formulation extends naturally to multiple batches. The forward pass of a single 2D convolutional layer [33] is
$Y_{n,h,w}=\sum_{c=1}^{C}\sum_{k=1}^{K}\sum_{l=1}^{L}X_{c,h+k,w+l}\ast\Theta_{n,c,k,l}+\beta_{n}$ (1)
where:
  • $X\in\mathbb{R}^{C\times H\times W}$: input tensor (batch size = 1)
  • $\Theta\in\mathbb{R}^{N\times C\times K\times L}$: filter tensor
  • $\beta\in\mathbb{R}^{N}$: bias vector
  • $n$: output channel index $(1\leq n\leq N)$
  • $h,w$: output spatial coordinates
  • $c$: input channel index $(1\leq c\leq C)$
  • $k,l$: kernel indices $(1\leq k\leq K,\;1\leq l\leq L)$
This convolution requires approximately $N\times H\times W\times C\times K\times L$ MAC operations per layer, and across layers, total MACs can reach tens of millions, imposing heavy demands on compute and memory bandwidth [13]. Define
$\tilde{Y}_{n,h,w}=\sum\limits_{c=1}^{C}\sum\limits_{k=1}^{K}\sum\limits_{l=1}^{L}X_{c,h+k,w+l}\ast\Theta_{n,c,k,l}$ (2)
To address this issue, the im2col transformation [7] enables tile-based GEMM, allowing a single MAC array to be reused across layers. Applied to (2), the transformation generally follows these steps:
  • For each output position $(h,w)$, a patch is extracted from the input:
    $\text{patch}_{h,w}(c,k,l)=X_{c,h+k,w+l},\quad\text{with shape }\mathbb{R}^{C\times K\times L}.$
  • Each extracted patch is flattened into a column vector:
    $\text{col}_{X}(h,w)\in\mathbb{R}^{CKL}\triangleq x_{i},\quad\text{with }N\triangleq CKL.$
  • Each filter $\Theta_{n}$ is flattened into a row vector:
    $\text{row}_{\Theta}(n)\in\mathbb{R}^{N}\triangleq\theta_{i}.$
  • The convolution is expressed as an inner product of length $CKL$ between the flattened input patch and filter:
    $\tilde{Y}=\sum_{i=1}^{CKL}\theta_{i}\,x_{i}+\beta_{i}.$ (3)
The resulting output is then reshaped into the appropriate order. For consistency in IPC formulation, we set $C=1$ and $L=1$ for simplicity, while the analysis can be extended for any $C$ and $L$. By denoting the summation in (3) as $\tilde{\tilde{Y}}$, we have
$\tilde{\tilde{Y}}=\sum_{i=1}^{K}\theta_{i}\,x_{i}.$ (4)
In (4), the IPC can be computed by representing either the inputs $x_{i}$ or the weights $\theta_{i}$ in OBC form [23]. For clarity, we refer to these as Scheme A and Scheme B, corresponding to the OBC representation of inputs $x_{i}$ and weights $\theta_{i}$, respectively. For Scheme A, each input $x_{i}$ with bit-width of $B_{1}$-bits is first represented in TC form as
$x_{i}=-\,x_{i,0}+\sum_{r=1}^{B_{1}-1}x_{i,r}\,2^{-r}$ (5)
where $b_{i,r}\in\{0,1\}$ is the $r$th bit of $x_{i}$. In OBC representation, $x_{i}$ can be expressed as $x_{i}=\frac{1}{2}[x_{i}-(-x_{i})]$, results (5) in
$x_{i}=\frac{1}{2}\left[\sum_{r=0}^{B_{1}-1}\Delta x_{i,r}2^{-r}-2^{-(B_{1}-1)}\right]$ (6)
with
$\Delta x_{i,r}=(-1)^{\lfloor r/(B_{1}-1)\rfloor}\left(x_{i,r}-\bar{x}_{i,r}\right)$ (7)
where $\bar{x}_{i,r}$ is the ones complement of $x_{i,r}$ and $\lfloor\cdot\rfloor$ is the floor function. Substituting (6) and (7) into (4) and rearranging, we have
$\tilde{\tilde{Y}}=\sum_{r=0}^{B_{1}-1}\Bigl(\sum_{i=1}^{K}\tfrac{1}{2}\,\theta_{i}\,\Delta x_{i,r}\Bigr)\,2^{-r}-\Bigl(\tfrac{1}{2}\sum_{i=1}^{K}\theta_{i}\Bigr)\,2^{-(B_{1}-1)}$ (8)
Define
$\tilde{\tilde{Y}}_{r}=\tfrac{1}{2}\sum_{i=1}^{K}\,\theta_{i}\,\Delta x_{i,r}\hskip 5.69046pt\textnormal{and},\hskip 5.69046pt\theta_{ot}=-\tfrac{1}{2}\sum_{i=1}^{K}\theta_{i}$ (9)
Substituting (9) into (8), we arrive at
$\tilde{\tilde{Y}}=\sum_{r=0}^{B_{1}-1}\tilde{\tilde{Y}}_{r}\,2^{-r}+\theta_{ot}\,2^{-(B_{1}-1)}$ (10)
Substituting (10) into (1) and ignoring the various subscripts, we get
$\tilde{Y}=\sum_{r=0}^{B_{1}-1}\tilde{\tilde{Y}}_{r}\,2^{-r}+(\theta_{ot}+\beta_{i}2^{B_{1}-1})2^{-(B_{1}-1)}$ (11)
where $\tilde{\theta}_{ot}=\theta_{ot}+\beta_{i}2^{B_{1}-1}$ indicates that an offset term is combined with a scaled bias term. Thus, (11) leads to
$\tilde{Y}=\sum_{r=0}^{B_{1}-1}\tilde{\tilde{Y}}_{r}\,2^{-r}+\tilde{\theta}_{ot}2^{-(B_{1}-1)}$ (12)
Each LUT entry corresponds to a possible combination of weights with input bit-slices $(x_{n,B_{1}-1},\ldots,x_{n,0};\ 1\leq n\leq K)$, ordered from least to most significant bit. During runtime, these bit-slices serve as addresses to access the LUT. For each address access, the output is accumulated using a SA operation, repeated over $B_{1}$ clock cycles to compute the output $\tilde{Y}$, as shown in Fig. 2(a). Upon closer observation, the offset term $\theta_{ot}$ corresponds to 2’s complement of the LUT content stored at the $0^{\textnormal{th}}$ address location. This value is accessed when loading the offset term at the beginning of shift-accumulation, as illustrated in Fig. 2(a). While traditional LUTs can store weight combinations, they become inefficient for large $K$ due to increased access time. An alternative in [26] uses adders and multiplexers to generate combinations dynamically but lacks scalability and introduces significant hardware overhead. Specifically, implementing such a LUT requires $2^{K}(1-2^{-K/2})$ adders/subtractors and $(2^{K-1}-1)$ 2-to-1 multiplexers. The CPD is $\log_{2}K\cdot T_{A}+(K-1)\cdot T_{MX}$, where $T_{A}$ and $T_{MX}$ are the delays of an adder and a multiplexer, respectively. To mitigate this, carry-save full adders (CSFAs) have been used to reduce delay.
Similar to (12), an expression for $Y$ can be obtained with Scheme B, where each weight $\theta_{i}$, represented in OBC form with a bit-width of $B_{2}$, follows the steps indicated by (5)–(12). In this case, we would have
$\tilde{Y}=\sum_{r=0}^{B_{2}-1}{\tilde{\tilde{Z}}}_{r}\,2^{-r}+{\tilde{x}}_{ot}2^{-(B_{2}-1)}$ (13)
where the variables in (13) are defined as follows:
$\tilde{\tilde{Z}}_{r}$ $=\tfrac{1}{2}\sum_{i=1}^{K}\,x_{i}\,\Delta\theta_{i,r}$
$\Delta\theta_{i,r}$ $=(-1)^{\left\lfloor r/(B_{2}-1)\right\rfloor}\left(\theta_{i,r}-\bar{\theta}_{i,r}\right)$
$\tilde{x}_{{ot}}$ $=x_{{ot}}+\beta_{i}\,2^{B_{2}-1},\hskip 5.69046pt\textnormal{and}\hskip 5.69046ptx_{{ot}}=-\tfrac{1}{2}\sum_{i=1}^{K}x_{i}.$
Note that Scheme B has the same structure as Scheme A (Fig. 2), except that the LUT content and the number of clock cycles in the SA unit may differ ($B_{1}\neq B_{2}$). These schemes differ only in choosing which component to represent in OBC form, exploiting the asymmetry between weight and input bit-widths to provide hardware implementation trade-offs.
III Proposed Architecture
III-A System Overview
The proposed CNN accelerator is implemented on the Xilinx ZCU106 platform. All input feature maps, weights, and biases are pre-loaded into on-chip RAM before execution, and the computed outputs are written back to the same memory. To favor edge deployment and design flexibility, all buffers are implemented using Slice RAM rather than dedicated BRAM. The accelerator design workflow consists of two stages: (1) model formulation with hardware-oriented optimization, and (2) deployment on the FPGA. The CNN model is first expressed using the two OBC schemes (A and B), which define the GEMM decomposition strategy, and is then mapped to hardware-optimized architectures.
[…]
Figure 3: System-level diagram of the proposed CNN accelerator.
III-B Address Generator and Control Unit
At the core of the address generator is a set of nested counters operating under the FSM’s supervision, responsible for address generation and read/write data scheduling. As shown in Fig. 4, the counter architecture consists of a dedicated cntr0 and three hierarchical counter groups for read (rd_cntr#$m$), calculation (cal_cntr#$m$), and write (wr_cntr#$m$) operations. Each group contains four levels of counters indexed by $m\in\{1,4\}$, corresponding respectively to tiling (i.e., partitioning a convolution into smaller blocks for efficient computation), spatial position $(h,w)$ as defined in (1), output channel index $(n)$, and layer index. The standalone cntr0 tracks the clock cycle index within the OBC-GEMM computation. The signals marked 1 to 4 in Fig. 4 represent the hierarchical carry mechanism, which enables stepwise progression within each counter group and governs the transfer of counter values between the read, calculation, and write stages.
In this hierarchical carry mechanism, hierarchical counter rippling applies primarily to the read counter stage. The clock cycle index counter (cntr0) generates sequential memory addresses for data fetches within one tile-level OBC-GEMM operation. When it reaches its upper bound (equal to the number of fetch cycles per tile), all data for the current tile has been loaded. At this point, carry signal 1 is asserted, cntr0 resets, and rd_cntr1 (tile index) increments to begin loading the next tile. When rd_cntr1 completes its range (number of tiles per spatial location), carry 2 is raised, rd_cntr1 resets, and rd_cntr2 (spatial position) increments. Similarly, once rd_cntr2 reaches its bound (number of spatial positions per output channel), carry 3 triggers, resetting rd_cntr2 and incrementing rd_cntr3 (output channel index). Finally, when rd_cntr3 reaches the total output channels for the layer, carry 4 asserts, resetting rd_cntr3 and incrementing rd_cntr4 (layer index) to start the next layer.
Figure 4: Hierarchical counter mechanism enabling address generation across read, calculate, and write stages.
[…]
III-C OBC-GEMM Core
The OBC-GEMM core consists of an array of $L$ LUTs of each size $K$, $L$ SA units, and $L$ PISO units of each size $K$ that handle bit-slices of inputs (or weights), as shown in Fig. 5. The LUTs dynamically generate partial sums using adders and multiplexers for hardware acceleration of IPCs, while SA units efficiently combine these results. PISO units bit-serialize the inputs (or weights) to fit the data flow of the scheme A (or scheme B). A key optimization is a single dedicated adder that computes the offset and bias terms. This eliminates separate bias computation, reducing hardware complexity, and saving area and power.
Figure 5: Top-level schematic of the proposed OBC-GEMM core.
III-C1 Parallel LUT
[…]
Figure 6: LUT architectures for $K=4$ based on the proposed techniques: (a) Parallel, (b) Shared, (c) Split, and (d) Hybrid.
III-C2 Shared LUT
As the name suggests, the shared LUT approach is derived from the parallel LUT by sharing the generated LUT contents to reduce the generated partial sums, as illustrated in Fig. 6(a). Specifically, it is observed that the LUT contents at address locations 4–7 can be derived by simply loading $\theta_{2}$ instead of $\theta_{2}$ through a 2-to-1 multiplexer, allowing a reduction in the computational cost. However, applying this optimization directly to the design in Fig. 2(b) is less straightforward due to the exponential growth in the number of adders across stages in the parallel architecture. By strategically sharing partial sums across mirrored address locations (e.g., 0–3 and 4–7), the proposed shared LUT design effectively reduces the number of required computations by nearly 50%.
III-C3 Split LUT
In both parallel and shared LUT techniques, the number of adders and multiplexers grows exponentially. Although shared LUTs halve the computation, they still exhibit significant exponential scaling. To alleviate this, the LUT is split into two parts. The first $P$ terms in the summation defined by $\tilde{Y}_{r}$ in (9) can be implemented using the parallel LUT, requiring $(2^{P-1}-1)+(P-1)$ adders and a $2^{P-1}$-to-1 multiplexer (or equivalently, $2^{P-1}-1$ two-to-one multiplexers). The remaining terms are computed similarly, using $(2^{K-P-1}-1)+(K-P-1)$ adders and a $2^{K-P-1}$-to-1 multiplexer. One extra adder combines the two partial results. Thus, the total number of adders is $(2^{P-1}+2^{K-P-1}-2)+(K-2)$, and the total number of 2-to-1 multiplexers is $2^{P-1}+2^{K-P-1}-2$. A key observation is that decreasing $P$ reduces complexity in the first part but increases it in the second, and vice versa. By minimizing the total hardware with respect to $P$, it can be shown that the optimal configuration occurs at $P=\frac{K}{2}$. This implies the most efficient design is achieved when both partitions are of equal size. The proposed LUT architecture based on this technique is shown in Fig. 6(c). It can be observed that the chain of adders is replaced by a single adder for the first $K/2$ section, while the other term is computed using a subtractor. The remaining two partial sums are simply the two’s complements of these results. A similar structure applies to the second $K/2$ section as well.
III-C4 Hybrid LUT
Although the split LUT can reduce the computational challenges associated with parallel and shared LUTs, it still exhibits exponential dependence, which may be undesirable for larger values of $K$. To overcome this limitation, a hybrid LUT is proposed, as shown in Fig. 6(d), which pairs consecutive weights and utilizes minimal logic resources for the generation of partial sums. The term hybrid here refers to the transformation of large-scale parallelism into a smaller degree of parallelism, inspired by the serial-like pairing of weights. In this approach, adders and subtractors are used to compute the partial sums $\theta_{K-i+1}+\theta_{K-i}$ and $\theta_{K-i+1}-\theta_{K-i}$ for all $i\in\{1,2,\dots,K-2\}$, and can be shared as they are independent of selection lines. The last pair of weights can be generated using either a conditional adder or an adder/subtractor, depending on the design requirements. In this case, a conditional adder is employed to minimize resource usage. As a result, each pair of weights requires only one adder or subtractor, resulting in a linear scaling of resource requirements, which is significantly more efficient compared to other techniques. Furthermore, the selection lines for the 2-to-1 multiplexers—used to select between $\theta_{K-i+1}+\theta_{K-i}$ and $\theta_{K-i+1}-\theta_{K-i}$—can be derived from the observation that the contents at the $0^{\textnormal{th}}$ and $3^{\textnormal{rd}}$ address locations are two’s complements of each other, as are the $1^{\textnormal{st}}$ and $2^{\textnormal{nd}}$, and similarly for the second half, as illustrated in Fig. 2(a).
TABLE I: Theoretical Hardware and Time Complexities
of the Proposed LUT Architectures with $K=p\times q$
Technique Adders 2-to-1 Muxes CPD
Parallel $(2^{q-1}+q-2)p+p-1$ $(2^{q-1}-1)p$ $(q+\textnormal{log}_{2}p)T_{A}$
Shared $(2^{q-2}+q-2)p+p-1$ $2^{q-2}p$ $(q+\textnormal{log}_{2}p)T_{A}$
Split $(2\cdot(2^{q/2-1}-1)+1)p+p-1$ $2\cdot(2^{q/2})p$ $(q/2+1+\textnormal{log}_{2}p)T_{A}+T_{MX}$
Hybrid qp+p1qp+p-1 3(q2)+13(q-2)+1 $(2+{\textnormal{log}}_{2}p)T_{A}+2T_{MX}$
  • Hybrid LUT has also qp/2qp/2–AND and qp/4qp/4–XOR gates to generate select signals for 2-to-1 multiplexers. $T_{A}$ and $T_{MX}$ are the computational delays of an adder and a 2-to-1 multiplexer respectively.
III-D Analysis of Co-Optimization Trade-offs in LUTs
III-D1 Hardware and Time Complexities
The theoretical estimates of hardware complexity (in terms of adders and 2-to-1 multiplexers) and time complexity (in terms of CPD) for the proposed LUT architectures are listed in Table I. These are estimated with the decomposition $K=p\times q$, to ensure fair comparison across designs. The Parallel LUT incurs the highest hardware cost due to exponential growth in both adders and multiplexers with respect to $q$. The Shared LUT reduces the adder count by reusing partial computations but retains similar CPD. The Split LUT lowers adder usage further, though this comes with an increase in multiplexers. Compared to the shared LUT, the Hybrid LUT offers the best balance, achieving linear adder scaling and reduced multiplexer count with respect to $q$. Notably, its CPD also remains low due to shallow adder trees. For a consistent comparison between different proposed LUT techniques, we set $q=4$, which implies $K=4p$. To validate these findings, all LUT architectures are implemented on FPGA with 8-bits symmetric operands for evaluation, as discussed in the following subsection.
Figure 7: LUT usage and power consumption versus clock frequency for different proposed techniques with $K=8,16,$ and $32$.
Figure 8: FF usage versus clock frequency for different techniques for $K=8,16,$ and $32$. The same markers are used as in Fig. 7.
III-D2 LUT and FF Usage vs Clock Frequency
Fig. 7 shows the resource utilization of the proposed LUT architectures implemented on FPGA across a clock frequency range of 100–800 MHz in 100 MHz steps. This evaluation examines how each design behaves as the clock frequency increases. LUT usage remains largely stable when timing constraints are easily met. At higher frequencies, however, FPGA synthesis and place-and-route tools introduce additional optimization logic—such as logic duplication and register insertion—to satisfy tighter timing requirements. This increase reflects tool-driven overhead rather than the intrinsic LUT requirements of each design. To ensure consistent comparisons, the frequency range is restricted. All designs exhibit LUT usage consistent with the complexities listed in Table I. The Parallel LUT shows the highest LUT usage due to its large number of adders and multiplexers. The Shared LUT reduces the LUT count by reusing partial computations, achieving moderate savings over Parallel LUT. The Split LUT reduces adders but raises logic utilization from added 2’s complementers; for $K=32$, performance is slightly worse due to the extra 2’s complement operations. The Hybrid LUT demonstrates the most efficient resource usage overall owing to optimized adders and minimal multiplexer overhead. Small fluctuations in LUT usage may occur across all designs due to synthesis heuristics and routing decisions, but these variations are minor.
Unlike LUT usage, flip-flop (FF) utilization remains largely consistent across all LUT designs, as shown in Fig. 8. This is because LUTs are primarily combinational, while FFs are used to store input/output feature maps and intermediate pipeline values, which are independent of the LUT architecture. When the kernel size $K$ increases, the datapath widens, requiring more FFs to hold intermediate results. This increase is similar across all designs, as $K$ affects sequential storage rather than the combinational structure of the LUTs.
III-D3 Power Consumption vs Clock Frequency
Accurate power estimation requires realistic activity data, as post-synthesis estimates often miss switching and routing effects in complex LUTs. In Xilinx Vivado, the process is: (1) run functional simulation; (2) generate a .saif file with signal toggles; (3) import it into power analysis; (4) combine toggles with utilization, timing, and clock frequency. This yields power figures that capture actual activity and frequency behavior. Using this procedure, power consumption increases with clock frequency for all LUT architectures due to higher switching activity and dynamic power. The slope of the power versus frequency curve varies across designs, reflecting differences in logic complexity, routing, and the number of active switching elements. The Parallel LUT exhibits the steepest slope because its large number of adders and multiplexers produces more toggling per cycle. The Shared LUT shows low power for $K=8$ because extensive computation reuse reduces switching activity, despite slightly higher combinational complexity. The Hybrid LUT, while consuming slightly more power than the Shared LUT at small $K$ due to active logic per patch, achieves more consistent power for $K=16$ and the lowest power at $K=32$ by efficiently pairing adders and minimizing multiplexers, thereby reducing toggling per operation. The Split LUT has shallower slopes than Parallel LUTs, as reduced adder counts lower overall switching activity. These trends show that power depends on both logic complexity and switching/data reuse.
TABLE II: Description of the Original and Modified LeNet‑5
Layer Original [34] Modified
Conv1 $5\times 5$ conv, 6 ch, tanh $5\times 5$ conv, 6 ch, ReLU
Pool1 $2\times 2$ avg pool, stride 2, tanh $3\times 3$ conv, stride 2, ReLU
Conv2 $5\times 5$ conv, 16 ch, tanh $5\times 5$ conv, 16 ch, ReLU
Pool2 $2\times 2$ avg pool, stride 2, tanh $3\times 3$ conv, stride 2, ReLU
GAP global avg pool on 16 ch
FC1 dense $400\!\to\!120$, tanh dense $16\!\to\!32$, ReLU
FC2 dense $120\!\to\!84$, tanh dense $32\!\to\!10$, softmax
FC3 dense $84\!\to\!10$, softmax
  • Channel counts preserved where applicable; FC layer dimensions are reduced to match modified spatial dimensions.
IV Model Considerations, Accuracy
and Evaluation of the Proposed OBC-GEMM
IV-A Model Considerations
To enable efficient FPGA implementation, we apply some lightweight modifications to the baseline LeNet-5 architecture [34], as listed in Table II:
  • The original use of tanh activations introduces saturation and requires scaling. Replacing them with ReLU not only yields higher sparsity but also enables simpler fixed-point implementation.
  • The original $2\times 2$ average-pooling layers neither learn parameters nor map cleanly onto the GEMM datapath. Replacing them with learnable $3\times 3$ strided convolutions unifies computation [35], increases arithmetic intensity, and better exposes the advantages of the proposed design.
  • LeNet-5’s two large FC layers dominate the parameter count. Substituting them with a GAP layer followed by a compact $32$-unit FC drastically reduces memory traffic while preserving accuracy [21].
  • One-side zero padding is inserted before each strided-convolution down-sampling step to keep spatial dimensions consistent, simplifying address generation and buffer reuse.
These modifications are adopted to demonstrate that the OBC-GEMM core applies cleanly to modern workloads. With fixed quantization, frequency, and tiling, they remain orthogonal to throughput. Specifically, replacing $2\times 2$ average pooling with $3\times 3$ stride-2 convolutions increases MACs per inference, while swapping tanh for ReLU and adding zero padding incur negligible MAC costs. GAP reduces parameters and memory traffic but not MACs per cycle. As in [20], these modifications mainly affect accuracy and memory; therefore, it is necessary to explore their impact in fixed-point implementations.
Figure 9: Accuracy of original and modified LeNet-5 under QAT with two sets of input bit-widths ($B_{1}$) and varying weight bit-widths ($B_{2}$).
IV-B Accuracy Comparison of LeNet-5 vs Modified LeNet-5
To evaluate the impact of lightweight architectural modifications under quantization-aware training (QAT), we compared the baseline LeNet-5 with a modified variant on MNIST, using padded $32\times 32$ images. Both models were trained with fixed-point arithmetic, fixing input bitwidth at $B_{1}=16$ or $B_{1}=8$, while varying weight bitwidth $B_{2}$. Results are summarized in Fig. 9.
For $B_{1}=16$ (top subplot), the baseline LeNet-5 achieved $91.96\%$$92.58\%$, with no consistent improvement as $B_{2}$ increased. In contrast, the modified network consistently outperformed the baseline, reaching $98.28\%$ at $B_{2}=4$ and $99.16\%$ at $B_{2}=16$. Gains were largest at mid-range bitwidths, where accuracy was nearly seven points higher. This shows that under higher input bitwidth, the modified architecture is far more resilient to weight quantization, sustaining near–floating-point performance even at low bit-widths.
With reduced input bitwidth ($B_{1}=8$, bottom subplot), the baseline degraded sharply at low $B_{2}$, dropping to $83.18\%$ at 2 bits and recovering to $94.34\%$ at 8 bits. The modified model was even more sensitive at $B_{2}=2$ ($64.32\%$), but rebounded strongly thereafter, achieving $98.22\%$ at $B_{2}=4$ and $98.60\%$ at $B_{2}=8$, exceeding the baseline by over four points at higher bitwidths.
Overall, two key insights emerge. First, replacing tanh with ReLU and introducing stride-2 convolutions significantly improve quantization robustness, particularly under moderate-to-high bitwidth. Second, while the modified LeNet-5 is vulnerable at ultra-low bitwidth ($B_{2}=2$ with $B_{1}=8$), performance recovers quickly with modest bit-width increases. These results confirm that careful architectural adjustments enable efficient hardware deployment, where reduced bitwidth yields memory savings, lower power use, and faster inference.
IV-C Evaluation of the OBC-GEMM Module
Figure 10: LUTs/$F$s usage and power of OBC-GEMM for different $(B_{1},B_{2})_{S}$ with $S\in\{A,B,AB\}$ indicates Scheme A or B or Symmetric case - AB values with hybrid and split LUT techniques at 50 MHz and 200 MHz.
For evaluation purposes, we discuss the FPGA implementation results corresponding to the accuracy outcomes for input bitwidth $B_{1}=16$ while varying the weight bitwidth $B_{2}$; however, the analysis could be further expanded to include $B_{1}=8$. The proposed OBC-GEMM module naturally supports operands with asymmetric quantization under Scheme A and Scheme B, as described in Section II. In this study, we fix the input bit-width at $B_{1}=16$ and vary the weight bit-width $B_{2}$ according to the accuracy results shown in Fig. 9. The case of $B_{2}=4$ with $B_{1}=16$ under Scheme A is omitted due to significant accuracy degradation, but the corresponding $(16,4)_{B}$ configuration is retained to highlight Scheme B’s behavior when weights are aggressively quantized.
Under asymmetric quantization, both proposed schemes follow $B_{1}>B_{2}$. In Scheme A, the SA unit operates for $B_{1}$ clock cycles corresponding to the input bits, while the partial products of the weights are generated using efficient hardware techniques, reducing arithmetic complexity and overall hardware resource usage. In contrast, Scheme B (Fig. 5) bit-slices the weights with bit-width $B_{2}$ through a PISO for address generation, while inputs in OBC form are used to generate the LUT data. Here, the data remain in the SA unit for $B_{2}$ clock cycles, enabling faster output, but the wider LUT data ($B_{1}$ bits) increase logic utilization and power consumption. The $(16,4)_{B}$ case illustrates this trade-off in its most extreme form: very short SA operation but significantly wider LUT paths. After integrating the Split and Hybrid LUTs and implementing both schemes, the results are summarized in Fig. 10. To evaluate the effect of timing constraints, we vary $B_{2}$, and $K$ at two operating frequencies: 50 MHz and 200 MHz. At 50 MHz, the designs meet timing with minimal replication, while at 200 MHz additional logic and routing are required to satisfy stricter constraints. Under both conditions, Scheme A and Scheme B exhibit consistent resource–performance trade-offs, and the choice of $(B_{1},B_{2})$ directly governs efficiency. The detailed effects on LUT usage, FF count, and power are discussed below.
IV-C1 LUT Usage vs. Clock Frequency
Fig. 10 reports the LUT utilization of the OBC-GEMM module with Split and Hybrid LUTs under different $(B_{1},B_{2})$ bitwidth pairs at 50 MHz and 200 MHz. Symmetric quantization occurs when $B_{1}=B_{2}$, i.e., $(8,8)_{AB}$ and $(16,16)_{AB}$, where the subscript $AB$ denotes the symmetric case with equal bitwidths for inputs and weights. Among these, $(16,16)_{AB}$ consumes the most LUTs, reflecting the high cost of fully 16-bit datapaths. Asymmetric input-dominant cases, such as $(16,12)_{A}$ and $(16,8)_{A}$, also exhibit substantial LUT demand. However, their usage remains lower than the symmetric baseline since the LUT width is limited by $B_{2}$. In contrast, Scheme B configurations, such as $(16,12)_{B}$, $(16,8)_{B}$, and $(16,4)_{B}$, incur higher LUT usage than their input counterparts, underscoring the stronger influence of weight precision in determining LUT cost compared to Scheme A. In fact, $(16,4)_{B}$ shows that pushing $B_{2}$ down to 4 substantially increases LUT pressure, since the LUT must still encode 16-bit input slices. Across all configurations, the Hybrid LUT design consistently requires fewer LUTs than the Split design. Hybrid packing achieves more efficient implementation by sharing PPs, while Split designs incur additional overhead from duplicated logic and reduced computation sharing. This efficiency gap becomes particularly pronounced at higher precision.
IV-C2 FF Usage vs. Clock Frequency
FF utilization, shown in Fig. 10, demonstrates a different trend compared to LUTs. Unlike LUT usage, FF demand is largely insensitive to clock frequency, remaining relatively stable between 50 MHz and 200 MHz. Instead, FF usage is primarily dictated by datapath width. Asymmetric configurations follow the expected scaling: pairs involving a 16-bit operand, such as $(16,12)_{A}$, $(16,8)_{A}$, $(16,12)_{B}$, $(16,8)_{B}$, and $(16,4)_{B}$, consistently fall between the symmetric and fully narrow cases. Notably, $(16,4)_{B}$ requires more FFs than its Scheme A counterparts of similar width, again reflecting the heavier storage overhead of LUT-based input expansion. Architecturally, the Split design consumes more FFs across all precision pairs, owing to the need for wider intermediate registers in its accumulation stage. In contrast, Hybrid LUT avoids unnecessary duplication and demonstrates greater efficiency, especially for mid-range precisions such as $(12,8)_{A}$, $(12,8)_{B}$ and $(8,8)_{AB}$.
IV-C3 Power vs. Clock Frequency
Power consumption, summarized in Fig. 10, increases noticeably with clock frequency across all configurations. The higher activity and deeper pipelining required to meet 200 MHz timing lead to significantly greater switching energy compared to 50 MHz. Furthermore, the precision scaling strongly impacts power consumption: the widest datapaths, including $(16,16)_{AB}$, $(16,12)_{A}$, and $(16,8)_{A}$, consistently incur the highest power consumption, while cases such as $(8,8)_{AB}$ are far more energy-efficient. Asymmetric configurations reveal an important tradeoff: e.g., $(16,8)_{A}$ requires more clock cycles but lower bits for generating the LUT contents in Scheme A, resulting in lower power compared to $(16,8)_{B}$, whereas in Scheme B, dominate switching activity despite shorter bit-width. The $(16,4)_{B}$ case amplifies this effect: the short cycle count reduces SA activity, but the extremely wide LUT data path drives high dynamic power, especially in the Split implementation. Between the two architectures, the Hybrid LUT design consistently consumes less power than the Split design. This difference is modest for shorter precisions but becomes substantial for wider datapaths. At $(16,16)_{AB}$ under 200 MHz operation, the Split architecture exhibits a steep increase in power due to redundant toggling in duplicated logic paths, while the Hybrid architecture maintains more controlled energy scaling.
Symmetric 16-bit quantization $(16,16)_{AB}$ is the most resource- and power-intensive, while narrower datapaths such as $(8,8)_{AB}$ improve energy and area efficiency. Asymmetric cases offer trade-offs: in Scheme A, logic and power consumption are reduced at the cost of requiring more cycles, whereas in Scheme B the situation is reversed. The inclusion of $(16,4)_{B}$ demonstrates the extreme of this trend—minimal cycles but substantially higher LUT and power cost—emphasizing the design space boundaries for OBC-GEMM. In all cases, the Hybrid architecture outperforms the Split architecture, saving LUTs, FFs, and power—especially at higher precision—thereby demonstrating the scalability of OBC-GEMM across bit-widths.
V Performance Comparison
V-A Performance Metric Definitions
The performance of ML accelerators is commonly expressed in terms of throughput, measured as operations per second (OP/s). Specifically, $\mathcal{T}_{\textnormal{MAC}}$ denotes the number of MAC operations per second. Since a MAC corresponds to a fused multiply–add, this convention is widely adopted for CNN workloads. From (3), throughput is expressed as
$\mathcal{T}_{\textnormal{MAC}}={\#\text{MACs}}/{T_{s}}={CKL}/{T_{s}},$ (14)
where $T_{s}$ is the sample period, assuming all MACs are available on hardware. Under the im2col transformation, these MACs can be equivalently expressed as $CL$ IPCs of size $K$:
$\mathcal{T}_{\textnormal{IPC}}={\#\text{IPC}}/{T_{s}}={CL}/{T_{s}}.$ (15)
In our work, all CNN computations are mapped onto the GEMM core; thus, CNN throughput is equivalent to GEMM throughput, expressed in IPCs as in (15). The proposed GEMM core instantiates only $L$ IPC blocks based on the OBC-DA (shown in Fig. 5), leading to the effective throughput
$\tilde{\mathcal{T}}_{\textnormal{IPC}}={\#{\text{IPC}}}/{CT_{s}}={L}/{T_{s}}.$ (16)
According to the OBC-DA principle, each IPC requires $B$ clock cycles for computation, independent of $K$ [26]. This is because partial products are generated while the SA unit produces outputs within $B$ cycles. Consequently, the sample rate $f_{s}$ must be scaled by $B$ (either $B_{1}$ for Scheme A or $B_{2}$ for Scheme B) relative to the clock frequency $f_{clk}$ to maintain synchronization in bit-serial processing. Using $T_{s}=1/f_{s}$ with $f_{s}=f_{clk}/B$, (16) simplifies to
$\tilde{\mathcal{T}}_{\textnormal{IPC}}=Lf_{s}=(L/B)f_{clk}.$ (17)
For fair comparison with MAC-based implementations, $\tilde{\mathcal{T}}_{\textnormal{IPC}}$ is translated to its MAC equivalent by multiplying with $K$, yielding
$\tilde{\mathcal{T}}_{\textnormal{MAC}}=({KL}/{B})f_{clk}.$ (18)
The equivalent number of slices (ENS) metric [36] provides a standardized measure to compare heterogeneous FPGA resources by converting them into a single, unified value. This allows fair and consistent evaluation across different FPGA families and designs. The ENS is defined as
$\textnormal{ENS}={\text{LUT}}/{4}+\text{DSP}\times 102.4+\text{BRAM}\times 116.2,$ (19)
[…]
The energy per sample (EPS) is an efficiency metric that captures the trade-off between power consumption and computational throughput. It is defined as
$\textnormal{EPS}={\text{Power}}/{\mathcal{T}_{\textnormal{MAC}}},$ (20)
By normalizing power against throughput, EPS quantifies the average energy required to process a single data sample. This measure enables fair comparison of different hardware implementations, as it reflects both computational performance and energy efficiency, which are critical for resource-constrained and high-performance FPGA applications.
Finally, to isolate architectural efficiency from clock frequency scaling, we normalize the throughput by both ENS and clock rate, as per
$\textnormal{AEP}={\mathcal{T}_{\textnormal{MAC}}}/({f_{clk}\times\textnormal{ENS}})$ (21)
where AEP denotes the Architectural Efficiency per ENS. This metric highlights the intrinsic efficiency of the architecture rather than gains achieved simply from higher clock frequency. By using ENS instead of DSP count, AEP captures the contribution of heterogeneous FPGA resources in a unified manner, ensuring that designs which rely heavily on BRAMs or LUTs are fairly compared against DSP-centric implementations. This normalization is particularly important when surveying prior work, where resource usage may vary widely, as it provides a balanced measure of architectural efficiency across different platforms and mapping strategies.
TABLE III: Performance Comparison of the State-of-the-Art CNN Accelerators Based on LeNet-5 Model on FPGA
[…]
Bitwidths 12/12 8/8 8/8 12/8 - 8/8 8/8 8/4 8/4 8/8 8/4 8/4
Model LeNet-5 Modified LeNet-5
Platform ZYBO Z7 ZCU106 XCVU9P Zynq-7020 Zynq-MPSoC Alpha Data 9H7 ZCU106
Freq (MHz) 100 150 300 100 100 250 100 100 95 100 100 95
LUTs 17420 514000 57657 18421 53292 16406 14724 23019 16310 14741 23071
$F$s 10195 28311 17472 1177 1044 1043 1170 1041 1041
DSPs 72 512 123 97 2614 0 0 0 0 0 0
BRAM 80 1,024 102 13.5 341 0 0 0 0 0 0
Power (W) 21.2 1.598 0.64 0.835 0.824 0.976 0.832 0.826 0.949
$\mathcal{T}_{\textnormal{MAC}}$ (GOP/s) 19.20 783.1 1490.5 0.141 3.02 110.8 0.2 0.2 0.38 0.2 0.2 0.38
ENS 21024 299918 38862 16107 320470 4102 3681 5755 4078 3685 5768
EPS 0.027 11.333 0.212 4.175 4.120 2.568 4.160 4.130 2.497
AEP 0.913 0.0166 0.0363 0.0018 0.001 0.488 0.543 0.695 0.490 0.543 0.694
  • ENS = LUT/4 + DSP $\times 102.4$ + BRAM $\times 116.2$ [36]. Assumptions: 1 DSP = $128\times 0.8=102.4$ slices, 1 BRAM (18k) = $166\times 0.7=116.2$ slices, 1 slice = 4 LUTs. EPS = Power / $\mathcal{T}_{\textnormal{MAC}}$ (W per GOP/s). AEP = $\mathcal{T}_{\textnormal{MAC}}/(f_{clk}\times\textnormal{ENS}/1000)$, reported in ops/cycle/kENS (kilo-ENS).
V-B Comparison with State-of-the-Art CNNs
[…]
  • LUT usage ranges from 14,724 to 23,071 and FFs from 1,041 to 1,177, while DSPs and BRAMs are entirely eliminated. This reduction in dedicated resources brings the ENS down to just 3,681–5,768 slices, far below DSP-centric designs.
  • GOP/s varies between 0.2–0.38. Some designs operate at a slightly lower frequency of 95 MHz due to timing closure constraints on the target FPGA, yet they maintain competitive throughput relative to their resource footprint.
  • EPS lies between 2.497–4.175 W/GOP/s, and AEP ranges from 0.488–0.695 ops/cycle/kENS. These values highlight highly efficient per-slice computation as well as favorable energy utilization compared to prior art.
  • In contrast to DSP- and BRAM-intensive approaches, the proposed designs show that high normalized efficiency can be achieved entirely with LUT-based computation, without the need for higher operating frequencies or large dedicated arithmetic resources.
The results clearly demonstrate that the proposed approach offers highly efficient resource utilization for low-precision CNN on resource-constrained platforms.
Table III highlights the contrast between prior DSP- and BRAM-intensive designs and the proposed DSP and BRAM-free approach. Conventional designs achieve high absolute throughput but require large FPGA resources:
  • [16] achieves the highest reported absolute throughput of 1,490.53 GOP/s using 512 DSPs, 1,024 BRAMs, 514,000 LUTs, and unspecified FFs at 300 MHz. However, the ENS is $\approx$299,918 slices, and the AEP is only 0.0166 ops/cycle/kENS, showing that this design prioritizes peak throughput at the expense of per-resource efficiency.
  • [15] reports 19.20 GOP/s using 72 DSPs, 80 BRAMs, 17,420 LUTs, and 10,195 FFs at 100 MHz. With an ENS of $\approx$21,024 and an AEP of 0.913 ops/cycle/kENS, this design exhibits better normalized efficiency than [16], but DSPs and BRAMs remain underutilized for low-precision operations.
  • [38] consumes 123 DSPs, 102 BRAMs, 57,657 LUTs, and 28,311 FFs for just 0.141 GOP/s at 100 MHz. The corresponding ENS is $\approx$38,862, EPS reaches 11.333 W/GOP/s, and AEP is only 0.0363 ops/cycle/kENS—clearly illustrating high resource and power overhead with limited throughput.
  • Other reported works [37, 39] achieve moderate throughput values but still depend heavily on DSPs and BRAMs, leading to low per-slice efficiency (AEP << 0.002) despite optimizations in arithmetic precision or resource mapping.
  • It is worth noting that [15] does not provide power consumption results, limiting a complete efficiency assessment. While its reported AEP appears higher than other works, the design occupies significantly more FFs than all the proposed variants, making it less favorable in terms of overall resource efficiency.
Overall, the proposed CNN accelerator achieves a favorable trade-off between throughput, FPGA resources (LUTs, FFs, DSPs, BRAMs), and energy efficiency (EPS, AEP). The Hybrid and Split variants along with Scheme A and B provide flexibility to tune enhance efficiency of the CNN inference. This confirms that the proposed designs are well-suited for low-power, resource-constrained edge AI applications.
VI Conclusion
In this work, efficient CNN implementations based on the proposed hardware OBC techniques have been presented. By formulating CNN inference in OBC form for both inputs and weights, bit-width asymmetry has been effectively exploited and its influence on hardware efficiency analyzed. Furthermore, OBC-based LUT implementations—parallel, shared, split, and hybrid—have been developed and evaluated to highlight their trade-offs in performance, resource usage, and energy efficiency. These optimizations have been integrated into an OBC-GEMM core using the im2col transformation, enabling streamlined and resource-aware CNN execution. These methods have been encapsulated within the COMET framework, providing a unified platform for co-optimized CNN acceleration. Experimental results on a fixed-point modified LeNet-5 have demonstrated substantial reductions in resource utilization compared to state-of-the-art CNNs based on LeNet-5 [16]. By leveraging FPGA LUT fabric and eliminating reliance on DSPs and BRAMs in the compute datapath, the proposed approach offers a scalable solution for lightweight CNN deployment on edge devices. Overall, this work establishes OBC-based computation as a promising and extensible paradigm for co-optimized CNN acceleration across diverse edge AI applications.